An urn contains four black balls and six white balls. What is the probability of getting one black and one white ball in two consecutive draws from the urn?

Answer by Boreal(15235)

BW or WB
Probability of first is 4/10*6/9=2/5 *2/3=4/15
probability of second is 6/10*4/9=3/5*4/9=4/15
we add them 8/15 is the answer.